OpenStudy (gabylovesyou):

LAST QUESTION FOR @jim_thompson5910 Assume a 125 pound person burns 150 calories after 15 min of jogging and 200 calories after 20 min of jogging. Part 1: Use calories as the y- coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the info provided?

OpenStudy (gabylovesyou):

Part 2: write an equation, in slope intercept form, to represent this data.

OpenStudy (gabylovesyou):

Part 3: How many calories will this person burn after 35 min of jogging? using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

First identify the points you are working with

jimthompson5910 (jim_thompson5910):

So one point for instance is (15, 150) because x = 15 and y = 150 (we're told that this person "150 calories after 15 min of jogging")

jimthompson5910 (jim_thompson5910):

What is the other point?

OpenStudy (gabylovesyou):

20,200

jimthompson5910 (jim_thompson5910):

now find the slope of the line through those two points

OpenStudy (gabylovesyou):

how :/

jimthompson5910 (jim_thompson5910):

subtract the y coordinates to get ??

OpenStudy (gabylovesyou):

200 - 150 = 50

jimthompson5910 (jim_thompson5910):

now subtract the x coordinates IN THE SAME ORDER

OpenStudy (gabylovesyou):

-5

jimthompson5910 (jim_thompson5910):

finally, divide the difference in y by the difference in x

jimthompson5910 (jim_thompson5910):

(50)/(-5) = ??

jimthompson5910 (jim_thompson5910):

oh wait, it should be +5 not -5

OpenStudy (gabylovesyou):

10

jimthompson5910 (jim_thompson5910):

so that's the slope

jimthompson5910 (jim_thompson5910):

what does it mean in this context?

OpenStudy (gabylovesyou):

ummm

OpenStudy (gabylovesyou):

how much he lost per minute

OpenStudy (gabylovesyou):

maybe lol

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (gabylovesyou):

ok can you put part 1 together? lol

jimthompson5910 (jim_thompson5910):

he burns 10 calories a minute

jimthompson5910 (jim_thompson5910):

wanna give it a shot and I'll tell you if you did it right or not?

OpenStudy (gabylovesyou):

to put it together?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (gabylovesyou):

(15,150) (20,200) slope:10 He burns 10 calories a min

jimthompson5910 (jim_thompson5910):

add in (200-150)/(20-15) = 50/5 = 10

OpenStudy (gabylovesyou):

before slope?

OpenStudy (gabylovesyou):

done...

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (gabylovesyou):

now part 2

jimthompson5910 (jim_thompson5910):

y = mx+b y = 10x+b 150 = 10(15) + b 150 = 150+b .. b = ??

OpenStudy (gabylovesyou):

0 ?

jimthompson5910 (jim_thompson5910):

so the equation is y = 10x

OpenStudy (gabylovesyou):

ok... got it! now part 3?

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

plug in x = 35 and tell me what you get

OpenStudy (gabylovesyou):

350

jimthompson5910 (jim_thompson5910):

so that's how many calories he burns in 35 min

OpenStudy (gabylovesyou):

ok how can the equation slope.... etc ??

jimthompson5910 (jim_thompson5910):

well you just used the equation to find how many calories he burned in 35 min

jimthompson5910 (jim_thompson5910):

when you graph the equation, you'll be able to see visually how many calories he burns at any given time

OpenStudy (gabylovesyou):

thank you sooo much!!! i really appreciate it!!!

OpenStudy (gabylovesyou):

your the best!! !!! im done with my test!!! WOOHOOOOO

jimthompson5910 (jim_thompson5910):

how did you do

OpenStudy (gabylovesyou):

all right :)

jimthompson5910 (jim_thompson5910):

that's great